### v2

# SAGE: Spatial-visual Adaptive Graph Exploration for Visual Place Recognition

Abstract
Visual Place Recognition (VPR) requires robust retrieval of geotagged images despite large appearance, viewpoint, and environmental variation. Prior methods focus on descriptor fine-tuning or fixed sampling strategies yet neglect the dynamic interplay between spatial context and visual similarity during training. We present SAGE (Spatial-visual Adaptive Graph Exploration), a unified training pipeline that enhances granular spatial–visual discrimination by jointly improving local feature aggregation, organize samples during training, and hard sample mining. We introduce a lightweight Soft Probing module that learns residual weights from training data for patch descriptors before bilinear aggregation, boosting distinctive local cues. During training we reconstruct an online geo–visual graph that fuses geographic proximity and current visual similarity so that candidate neighborhoods reflect the evolving embedding landscape. To concentrate learning on the most informative place neighborhoods, we seed clusters from high-affinity anchors and iteratively expand them with a greedy weighted clique expansion sampler. Implemented with a frozen DINOv2 backbone and parameter-efficient fine-tuning, SAGE achieves SOTA across eight benchmarks. It attains 98.9%, 95.8%, 94.5%, and 96.0% Recall@1 on SPED, Pitts30k-test, MSLS-val, and Nordland, respectively. Notably, our method obtains 100% Recall@10 on SPED only using 4096D global descriptors. Code and models will be released upon acceptance.
1 Introduction
Visual Place Recognition (VPR) matches a query image to its corresponding location within a large-scale geotagged database, serving as a fundamental capability for critical applications such as autonomous robot navigation (li2024navigation), loop closure detection for autonomous driving (Bench2drive), and large-scale map construction (ho2024map). The main challenge of VPR is maintaining robust retrieval performance under severe and unconstrained environmental changes, including extreme viewpoint shifts, illumination variations, adverse weather, long-term temporal drift, and frequent dynamic occluders, among others (tzachor2024effovpr; wang2025feature).
Early VPR methods relied on hand-crafted local descriptors (SIFT; SURF) and aggregated them into global encodings via pooling schemes such as Bag of Words or VLAD (BoW; vlad). However, these methods lack adaptability and perform poorly under large-scale appearance changes. With the advent of deep learning, learnable aggregation modules (netvlad; gem) were introduced, which enhanced descriptor compactness and robustness by learning task-specific pooling strategies. Subsequent research has mainly focused on simplifying, regularizing, or refining aggregation mechanisms to improve generalization and computational efficiency (jin2025edtformer), such as reducing reliance on explicit cluster centers or alleviating the “burstiness” problem of local features (supervlad; VLAD-BuFF).
Figure 1: Performance and parameter efficiency of SAGE. (a–d) Recall@1 across four datasets at different global descriptor dimensions; SAGE achieves the best performance regardless of backbone and descriptor size. (e) Parameter comparison. By freezing DINOv2, SAGE substantially reduces trainable parameters compared to methods employing adapters or partial encoder tuning methods, demonstrating high efficiency. (f) Recall@1 performance compared with EMVP across the datasets.
Recently, the advent of Visual Foundation Models (VFMs) (vit; dinov2) has advanced VPR by enabling the capture of long-range semantic dependencies and richer interactions between image patches (boq; FoL). These strategies reduce sensitivity to occlusion and background clutter and improve robustness with controlled parameter overhead. While recent adaptation methods for VFMs are notably efficient, different strategies within the broader VFMs ecosystem vary in resource consumption (EDTformer). For instance, fine-tuning a backbone’s encoder layers typically requires more computational resources than Parameter-Efficient Fine-Tuning (PEFT) approaches (jia2022visual; emvp). Furthermore, While the model’s understanding is dynamic, a pre-defined or static sampling policy (salad-cm) may fail to consistently present the most informative examples as learning progresses.
Recent research has explored constructing training batches that reflect real-world difficulty. However, a common strategy operates on a static “think-once, act-always” principle, relying on offline computations like pre-defined clustering based on initial features (QAConv-GS; GCL; salad-cm). This approach overlooks a crucial fact: what constitutes a hard sample is not a fixed property but a dynamic state that evolves together with the model’s embedding geometry during training. Effective mining thus requires an architectural “slow thinking” involving an iterative reassessment of difficulty. Without this, static methods quickly become obsolete. They continue feeding the model stale examples as old challenges turn trivial and new ones emerge at the decision boundary. This mismatch between a static sampling strategy and a dynamic learning process creates a critical bottleneck, hindering the model’s full discriminative potential.
To address these interconnected limitations, we propose SAGE (Spatial-Visual Adaptive Graph Exploration), a unified VPR training framework that embraces a “slow thinking” paradigm for hard sample mining. Rather than depending on a one-time, fixed policy that labels samples as hard for the entire training run, SAGE continuously revisits and updates the hardness labels in response to changes in the model’s representation. This philosophy is realized through a fundamentally dynamic architecture. At its core, an online process reconstructs a geo-visual affinity graph each epoch, ensuring the sampling strategy stays synchronized with the model’s evolving embedding space. To maximize the impact of this intelligent sampling, SAGE also incorporates lightweight modules, including Soft Probing (SoftP) and an InteractHead, which enhance descriptor quality by amplifying discriminative local patches and modeling cross-image associations. This synergy between dynamic mining and enhanced feature representation allows SAGE to focus learning on the most informative spatial-visual neighborhoods, leading to state-of-the-art accuracy with remarkable parameter efficiency, as summarized in Fig. 1. In summary, the main contributions of this paper are:
  • SoftP Feature Interaction. We propose SoftP, a lightweight module that uses data-driven residual weighting to enhance discriminative local patches, and an InteractHead that models associations between fragments across images, thereby improving descriptor coherence across views.
  • Dynamic Geo-Visual Graph Mining. Our online strategy dynamically rebuilds the geo-visual affinity graph each epoch, keeping the mining process aligned with the model’s evolving embedding space while prioritizing the most informative samples for faster convergence.
  • Weighted Greedy Clique Expansion. Our weight-guided algorithm initiates sampling from anchors with high affinity and iteratively expands the most challenging neighborhoods, thereby generating balanced batches of utility that focus learning on detailed spatial and visual distinctions.
  • Efficient SOTA Accuracy. Implemented with a frozen DINOv2 backbone and parameter-efficient fine-tuning, SAGE sets SOTA on eight VPR benchmarks (Fig. 1), retaining competitiveness even with compact descriptors.
2 Related Work
Visual Place Recognition (VPR) requires global descriptors to remain compact and robust under substantial variations in viewpoint, illumination, and scene structure. Early methods for generating global descriptors, such as NetVLAD (netvlad), utilized a vast set of cluster centers, which rendered them vulnerable to domain shifts. Although self-supervised hard sample mining (sfrs) has been introduced to mitigate this issue, such methods remain constrained by the representational capacity of CNN backbones, yielding suboptimal performance. With the incorporation of spatial attention (delf) and feature reweighting (solar), descriptor robustness has been yet improved. Other works focus on optimizing feature aggregation (mixvpr; salad; boq), which improves performance but relies on fixed aggregation strategies and thus lacks adaptability to dynamically evolving embeddings. Recent studies have highlighted the effectiveness of modulating feature magnitudes in a lightweight, data-dependent manner prior to their aggregation. Such approaches include non-local attention for adaptive spatial weighting (chen2023self), efficient context encoding (huang2022efficient), parameter-efficient tuning with second-order moments (gao2023tuning), and the exponentially weighted fusion of pooling kernels (stergiou2022adapool). Other methods reduce overhead by lowering the number of clusters or entirely eliminating clustering through centroid-free probes (CFP) (supervlad; emvp), yielding compact descriptors by utilizing second-order feature statistics. Two-stage approaches (selavpr; FoL) improve retrieval accuracy through local feature re-ranking but introduce additional computational overhead. Recent single-stage methods can achieve comparable or even superior performance using only global features (megaloc; superplace). This pursuit of efficiency is also reflected in adapting powerful Visual Foundation Models (VFMs) for downstream tasks (zhang2025d; embodiedplace). With the proliferation of VFMs, Parameter-Efficient Fine-Tuning (PEFT) has emerged as a crucial paradigm. Instead of fine-tuning the entire backbone, these methods update only a small subset of parameters, such as lightweight adapters or normalization layers, thereby significantly enhancing training efficiency (jia2022visual; emvp). Our method follows this paradigm by freezing the backbone while introducing lightweight modules to enhance feature discriminability. Cross-image correlation methods (cricavpr; emvp) further enhance matching by capturing and modeling inter-image dependencies. In line with this paradigm, our approach strategically enhances feature discriminability. We introduce Soft Probe, a lightweight residual module that adaptively amplifies salient local regions, and InteractHead, which models cross-image dependencies. Together, they significantly boost descriptor quality with minimal parameter overhead.
In deep metric learning, dynamic sampling strategies (DynamicSampling) adjust the importance of training pairs with epoch-dependent weighting terms, organizing them in an “easy-to-hard” order, which enables the network to first learn general category boundaries from easy samples and then focus on hard samples in later stages. More recently, a graph-based sampling method (QAConv-GS) has been proposed, which constructs a nearest neighbor graph from class embeddings at the beginning of each epoch. By selecting an anchor class and its neighboring classes to form training batches, this approach improves the discriminative power of learned embeddings and enhances training efficiency. In VPR, spatial graphs have been used to encode geographic relationships. For example, the MMS-VPR benchmark (MMSVPR) represents street intersections and road segments as nodes and edges, leveraging topological context to improve retrieval performance. Such approaches often rely on mining discriminative regions or focusing on hard positive samples (aanet; seidenschwarz2021learning; fang2022adversarial; FoL) to enhance accuracy and robustness, but ignore geo-information. Other works reformulate VPR as a classification task (cosplace; eigenplaces) to avoid explicit mining, but these remain limited by static sample selection and the neglect of geographic information. Moreover, on sparse datasets such as GSV-Cities (gsv), these limitations restrict generalization capability. On dense datasets such as MSLS (msls), offline clustering methods (salad-cm) partition visually similar and geographically neighboring images into fixed clusters for training. Another category of approaches leverages hybrid strategies (kalantidis2020hard), hard negative mining (garg2022seqmatchnet; deuser2023sample4geo; ali2023global) or generation (peng2024globally) to enhance retrieval performance. However, most of these approaches struggle to generalize effectively. In contrast to such static or scheduled strategies, SAGE reconstructs a geo-visual graph at every training epoch and employs greedy sampling to focus on the most densely populated and challenging clusters in the evolving embedding spaces, achieving superior performance across datasets.
3 Method
Fig. 2 illustrates the proposed SAGE framework. First, the frozen DINOv2 feature extraction backbone processes input images (Sec.3.1). Next, the Soft Probing (SoftP) module aggregates these features into robust global descriptors (Sec.3.2). Then, Online Graph Creation employs InteractHead to refine descriptors and integrates geographic and visual distances to form dynamic graphs (Sec.3.3). Finally, we adopt greedy weighted sampling to focus training on hard examples (Sec.3.4).
3.1 Feature Extraction
DINOv2 provides strong visual representations from large-scale self-supervised pretraining. We use a pretrained DINOv2 as a frozen backbone and achieve parameter-efficient fine-tuning by inserting learnable Dynamic Power Normalization (DPN) layers into the last $N$ encoder blocks (dinov2; emvp). As shown in Fig. 2(a), an input image $I\in\mathbb{R}^{H\times W\times 3}$ passes through two stages. Uncalibrated Blocks extract base features, and Recalibrated Blocks integrate DPN to recalibrate features and produce task-specific representations. The backbone outputs one learnable class token and $L$ patch tokens, forming a token matrix $\mathbf{f}\in\mathbb{R}^{(L+1)\times M}$, where $M$ is the embedding dimension. SoftP and InteractHead then aggregate and refine these tokens to strengthen cross-image correspondence and yield a discriminative global descriptor $\mathbf{F}\in\mathbb{R}^{D\times K}$ for place recognition.
Figure 2: SAGE overview. (a) Pipeline: a frozen DINOv2 with PEFT outputs tokens; SoftP amplifies informative patches, and InteractHead applies cross-image attention to form a robust global descriptor. (b) Online Graph Creation: each epoch builds a geo–visual affinity graph, keeping top-k neighbors and updating edges as embeddings evolve. (c) Greedy Weighted Sampling: seed by average affinity and expand cliques by adding the most connected nodes. (d) SoftP: A lightweight module that uses residual weighting to emphasize discriminative features prior to aggregation.
3.2 Soft Probing
Centroid-Free Probing (CFP) crafts a global image descriptor $\mathbf{f}$ by aggregating the $L$ local spatial descriptors $\{X_{i}\}_{i=1}^{L}$ into a second-order Gram-like covariance matrix, which eliminates the need for offline cluster centers (emvp; supervlad). Although related methods like Moment Probing also leverage second-order statistics (wang2022dropcov; gao2023tuning), a key limitation of CFP is its uniform treatment of all descriptors. This equal weighting can underemphasize subtle yet discriminative local cues. To address this shortcoming, we introduce Soft Probing (SoftP), a lightweight module shown in Fig. 2(d), which adaptively emphasizes informative spatial locations while preserving the underlying feature geometry before the final aggregation step.
SoftP first computes a scalar response for each descriptor and converts it into a bounded residual coefficient; this coefficient is broadcast across channels and applied in residual form to obtain the modulated descriptors. Concretely, for each descriptor $X_{i}$ we compute an $\ell_{2}$ response $s_{i}=\|X_{i}\|_{2}+\varepsilon$ (with $\varepsilon>0$ for numerical stability), and feed $s_{i}$ into a compact predictor $\phi$ (a two-layer MLP) that outputs a scalar which is squashed by a sigmoid and scaled by a hyperparameter $\alpha$:
$\beta_{i}\;=\;\alpha\cdot\sigma\!\bigl(\phi(s_{i})\bigr),\qquad 0\leq\beta_{i}\leq\alpha.$ (1)
The modulated descriptor is formed residually as:
$\widetilde{X}_{i}\;=\;X_{i}+\beta_{i}X_{i}\;=\;(1+\beta_{i})X_{i}.$ (2)
This residual reweighting behaves like a soft, data-driven attention mechanism: it amplifies salient responses while avoiding destructive rescaling of channel structure (solar). Under mild assumptions (or to first order when $\beta_{i}$ is small and the mean shift is negligible) (gao2023tuning), the variance of each dimension of the modulated descriptors increases approximately as:
$\mathrm{Var}(\{\widetilde{X}_{i}\})=\frac{1}{N}\sum_{i=1}^{N}(1+2\beta_{i})\,\bigl\|X_{i}-\bar{X}\bigr\|^{2}=\underbrace{\frac{1}{N}\sum_{i=1}^{N}\bigl\|X_{i}-\bar{X}\bigr\|^{2}}_{\mathrm{Var}(\{X_{i}\})}+\frac{2}{N}\sum_{i=1}^{N}\beta_{i}\,\bigl\|X_{i}-\bar{X}\bigr\|^{2}.$ (3)
where $\bar{X}=\tfrac{1}{N}\sum_{i}X_{i}$. This relation highlights that SoftP selectively enlarges the variance contribution of high-response locations, thereby enhancing the sensitivity of subsequent aggregation stages to discriminative local structures. Finally, the set of modulated descriptors $\{\widetilde{X}_{i}\}$ is passed to the aggregation stage to produce the final global descriptor. SoftP adds only a negligible number of parameters, preserves the semantic geometry of the original descriptors, and consistently improves the robustness of the resulting global descriptor under significant appearance changes.
3.3 Online Graph Creation
To prepare for the graph creation, the image descriptors $\mathbf{f}_{i}\in\mathbb{R}^{D\times K}$ produced by SoftP are first processed by the InteractHead. Departing from prior method that partition features via learned cluster assignments(cricavpr; supervlad), InteractHead deterministically splits each descriptor into $S$ fixed length segments, $\{\mathbf{f}_{i}^{(s)}\}_{s=1}^{S}$. To enable cross-image attention, these segments are rearranged such that for each index $s$, the segments from all $B$ images form a sequence. These sequences are then processed by a two-layer Transformer encoder ($\mathcal{E}$) with GELU activation. This encoder structure applies attention across the batch for each segment type, capturing consistent correlations across views and improving descriptor robustness. The enhanced descriptors $\mathbf{F}$ are obtained by:
$\mathbf{F}=\mathrm{reshape}\Bigl(\mathcal{E}\bigl([\mathbf{f}_{1}^{(1)},\dots,\mathbf{f}_{B}^{(1)};\dots;\mathbf{f}_{1}^{(S)},\dots,\mathbf{f}_{B}^{(S)}]\bigr)\Bigr),$ (4)
Our approach continuously aligns candidate graphs to the model’s current embedding space by reconstructing the graph at each training epoch. First, for each city we group images by their unique cluster labels and randomly sample one image from each cluster. The sampled images are passed through our model to obtain descriptors that serve as representative features for the clusters, producing a set of cluster-level features for every city.
The process begins by sampling cities with a probability proportional to their cluster count. From a chosen city, we randomly select a single cluster, termed a “place”. We then identify $P$ similar places by computing cosine distances between the descriptor of our selected place and all other cluster descriptors, and sampling probabilistically such that smaller distances yield a higher selection probability. The images from these $P+1$ total places become the unordered nodes of a graph, for which we compute all pairwise Euclidean geographic distances $d_{\mathrm{geo}}(i,j)$. Subsequently, we construct an adjacency graph by connecting nodes whose geographic distance is below a threshold $\tau$. From this graph, we extract several cliques (i.e., complete subgraphs), denoted as $G=(V,E)$. Finally, within each clique, we calculate the pairwise visual descriptor distances as $d_{\mathrm{vis}}(i,j)=\|\mathbf{F}_{i}-\mathbf{F}_{j}\|_{2}$. To combine these two distance types and do so multiplicatively, we define:
$W_{ij}=-\left(d_{\mathrm{geo}}(i,j)\cdot d_{\mathrm{vis}}(i,j)\right),\quad W_{ii}=0.$ (5)
As illustrated in Fig. 2(b), we construct a sparse affinity graph $\mathcal{G}=(V,E^{\prime})$, where an edge $(i,j)$ exists if its affinity score $W_{ij}$ exceeds a predefined threshold $\tau_{2}$. This graph is dynamic, as it is rebuilt each epoch to reflect the continuous evolution of the model’s embeddings. From this graph, we seek a complete subgraph, known as a clique, for the sampling process. The search concludes upon finding the first clique $C$ that meets a minimum size requirement of $|V_{C}|\geq N=10$. This clique is then used to guide the subsequent sampling stage.
3.4 Greedy Weighted Sampling
We propose a greedy, weight-driven selection process that adaptively focuses training on the most informative neighborhoods while fully leveraging the reconstructed geo-visual graph. We first identify the most central node in the graph to serve as a cluster anchor. This is achieved by computing a seed score $S(i)$ for each node, which represents its total affinity to all other nodes:
$S(i)=\frac{1}{N-1}\sum_{\begin{subarray}{c}j=0\\ j\neq i\end{subarray}}^{N-1}W_{ij}.$ (6)
Fig. 2(c) shows the node with the highest score is chosen as the initial member of our training clique, $C=\{v_{0}^{*}\}$, where $v_{0}^{*}=\arg\max_{i}S(i)$. Subsequently, we iteratively expand the clique $C$ by adding the node $v^{*}$ that exhibits the highest average affinity to the current members of $C$:
$v^{*}=\arg\max_{\begin{subarray}{c}v\notin C\end{subarray}}\frac{1}{|C|}\sum_{u\in C}W_{u,v}.$ (7)
This procedure is repeated until the clique reaches the desired size $|C|=k$, where $k$ = 4. By seeding from a central anchor and greedily expanding towards the closest nodes, our method effectively drills down into the densest subgraphs of the geo-visual landscape. These dense regions represent clusters of mutually confusing samples the most difficult scenarios where the model struggles to make fine-grained distinctions. Our approach is inherently adaptive to the model’s learning progress. It dynamically responds to the evolving weight distribution each epoch, concentrating training effort on the most pertinent hard positive and negative examples. This adaptive focus not only accelerates convergence but also enhances the model’s robustness against subtle spatial and visual ambiguities.
4 Experiments
4.1 Datasets and Performance Evaluation
We validate SAGE on a diverse collection of VPR benchmarks (Tab. 1) covering common real-world challenges: Pitts30k-test and Pitts250k-test (large viewpoint variation) (pitts), SPED (low-quality / high scene depth and condition changes) (SPED), MSLS-val (multi-year urban/suburban variability) (msls), Nordland (four-season natural scenes) (nordland), Tokyo247 (multi-view urban captures) (tokyo247), AmsterTime (historical grayscale vs. contemporary RGB) (amstertime), and Eynsham (rural grayscale route) (eynsham; berton2022deep). Further details can be found in App. A.3.
The experiment adopts Recall@N (R@N) as the evaluation metric, i.e., the percentage of query images for which at least one of the top-N retrieved database images geographically matches the query image (within a preset threshold). Thresholds follow standard protocols: 25 meters for Pitts30k-test, Pitts250k-test, Tokyo24/7, Eynsham and SPED; MSLS-val uses 25 meters with azimuth within 40 degrees; Nordland uses $\pm 10$ frames (pitts; SPED; msls).
Table 1: Summary of the evaluation datasets.
Dataset Description Number
Database Queries
Pitts30k-test urban, panorama 10,000 6,816
MSLS-val urban, suburban 18,871 740
Nordland natural, seasonal 27,592 27,592
SPED various scenes 607 607
Tokyo24/7 urban, time-varying 75,984 315
AmsterTime urban, time-related 1,231 1,231
Pitts250k-test urban, panorama 83,952 8,280
Eynsham rural, historical 23,935 23,935
Table 2: Comparison to SoTA Methods on VPR Benchmark Datasets. The best and second best metrics are shown in red bold and blue bold, respectively. Two-stage methods are denoted by .
Method Dim SPED Pitts30k-test MSLS-val Nordland
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
NetVLAD $_{{\color[rgb]{0,0,1}\text{CVPR' 2016}}}$ 32768 70.2 84.5 89.5 81.9 91.2 93.7 53.1 66.5 71.1 6.4 10.1 12.5
SFRS $_{{\color[rgb]{0,0,1}\text{ECCV' 2020}}}$ 4096 80.2 92.6 95.4 89.4 94.7 95.9 69.2 80.3 83.1 16.1 23.9 28.4
CosPlace $_{{\color[rgb]{0,0,1}\text{CVPR' 2022}}}$ 512 75.5 87.0 89.6 88.4 94.5 95.7 82.8 89.7 92.0 58.5 73.7 79.4
MixVPR $_{{\color[rgb]{0,0,1}\text{WACV' 2023}}}$ 4096 84.7 92.3 94.4 91.5 95.5 96.3 88.0 92.7 94.6 76.2 86.9 90.3
R2Former $_{{\color[rgb]{0,0,1}\text{CVPR' 2023}}}$ / 67.5 75.8 77.8 91.1 95.2 96.3 89.7 95.0 96.2 77.0 89.0 91.9
EigenPlaces $_{{\color[rgb]{0,0,1}\text{ICCV' 2023}}}$ 2048 70.2 83.5 87.5 92.5 96.8 97.6 89.1 93.8 95.0 71.2 83.8 88.1
SelaVPR $_{{\color[rgb]{0,0,1}\text{ICLR' 2024}}}$ 1024 83.5 92.6 94.6 90.2 96.1 97.1 87.7 95.8 96.6 72.3 89.4 94.4
SelaVPR $_{{\color[rgb]{0,0,1}\text{ICLR' 2024}}}$ / 88.6 95.1 97.2 92.8 96.8 97.7 90.8 96.4 97.2 87.3 93.8 95.6
CricaVPR $_{{\color[rgb]{0,0,1}\text{CVPR' 2024}}}$ 4096 91.3 95.2 96.2 94.9 97.3 98.2 90.0 95.4 96.4 90.7 96.3 97.6
SALAD $_{{\color[rgb]{0,0,1}\text{CVPR' 2024}}}$ 8448 92.1 96.2 96.5 92.5 96.4 97.5 92.2 96.4 97.0 89.7 95.5 97.0
EDTformer $_{{\color[rgb]{0,0,1}\text{TCSVT' 2025}}}$ 4096 92.4 95.9 96.9 93.4 97.0 97.9 92.0 96.6 97.2 88.3 95.3 97.0
BoQ $_{{\color[rgb]{0,0,1}\text{CVPR' 2024}}}$ 12288 92.5 95.9 96.7 93.7 97.1 97.9 93.8 96.8 97.0 90.6 96.0 97.5
SALAD-CM $_{{\color[rgb]{0,0,1}\text{ECCV' 2024}}}$ 8448 89.5 94.9 96.1 92.6 96.8 97.8 94.2 97.2 97.4 95.6 98.6 99.1
SuperVLAD $_{{\color[rgb]{0,0,1}\text{NIPS' 2024}}}$ 3072 93.2 97.0 98.0 95.0 97.4 98.2 92.2 96.6 97.4 91.0 96.4 97.7
EMVP $_{{\color[rgb]{0,0,1}\text{NIPS' 2024}}}$ 8448 94.6 97.5 98.4 94.0 97.5 98.2 93.9 97.3 97.6 88.7 97.3 99.3
FoL $_{{\color[rgb]{0,0,1}\text{AAAI' 2025}}}$ 8448 92.1 96.5 98.0 93.9 97.2 98.1 93.1 96.9 97.4 87.8 94.5 96.4
FoL $_{{\color[rgb]{0,0,1}\text{AAAI' 2025}}}$ / 92.6 96.5 97.4 94.5 97.4 98.2 93.5 96.9 97.6 92.6 96.7 97.8
2048 95.6 99.2 99.7 95.4 97.4 97.9 92.6 96.9 97.7 91.2 96.6 97.8
SAGE (Ours) 4096 97.7 99.8 100 95.6 97.7 98.3 93.7 97.3 97.8 94.4 98.2 99.0
8448 98.9 99.7 100 95.8 97.8 98.4 94.5 97.4 97.8 96.0 98.9 99.4
Table 3: Comparison to SoTA methods on more challenging datasets. Values marked with were reproduced in this work when they were not reported in the original publications.
Method Dim AmsterTime Tokyo24/7 Pitts250k‐test Eynsham
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
SALAD-CM $_{{\color[rgb]{0,0,1}\text{ECCV' 2024}}}$ 8448 57.8 77.5 81.3 96.8 97.5 97.8 95.2 98.8 99.3 91.9 95.3 96.1
SuperVLAD $_{{\color[rgb]{0,0,1}\text{NIPS' 2024}}}$ 3072 63.9 83.9 87.3 95.6 97.8 98.1 97.3 99.4 99.7 92.1 95.6 96.4
EMVP $_{{\color[rgb]{0,0,1}\text{NIPS' 2024}}}$ 8448 65.6 86.0 90.5 96.8 98.1 98.7 96.5 99.1 99.5 91.9 95.7 96.6
FoL $_{{\color[rgb]{0,0,1}\text{AAAI' 2025}}}$ 8448 64.6 84.3 88.2 96.2 98.7 98.7 96.5 99.1 99.5 91.7 95.3 96.2
2048 66.2 78.6 85.0 95.6 96.5 98.1 97.7 99.1 99.3 92.7 95.8 96.5
SAGE (Ours) 4096 76.0 88.0 92.3 96.5 99.1 99.4 98.2 99.4 99.5 92.9 96.0 96.8
8448 83.5 93.3 95.4 97.5 99.1 99.4 98.4 99.4 99.7 93.1 96.2 97.0
4.2 Implementation Details
SAGE is built upon the EMVP framework (emvp), which we reproduced from its publication for a fair comparison as the official code is unavailable. We fine-tune two Vision Transformer backbones, ViT-B and ViT-L, which we denote as SAGE-B and SAGE-L, respectively. The Feature Compression ($\mathcal{F}_{C}$) and Feature Probing ($\mathcal{F}_{P}$) branches are implemented as two-layer MLPs, reducing the feature dimensions to $D=128$ and $K=64$, respectively. The InteractHead module is implemented as a two-layer Transformer encoder with a model dimension of 768, 16 attention heads, and a feed forward network with dimension 1024. During training we freeze the backbone and adopt DPN for PEFT, which adaptively preserves task-specific information while greatly reducing the number of trainable parameters. Each training mini-batch is constructed with equal contributions from MSLS (all non-panoramic images from the training set) and GSV-Cities (0.56M images from 67K places). For every batch we build a new sparse geo-visual graph, and we sample $P=15$ sequences from the same city and recompute cliques dynamically during training. The thresholds for this process are set to $\tau_{1}=25$ m and $\tau_{2}=-2.88\times 10^{3}$. Input images are resized to $224\times 224$ during training and $322\times 322$ during inference. All experiments are implemented in PyTorch and run efficiently on a single NVIDIA A100 GPU. We fine-tune models for 10 epochs and select the checkpoint with the highest Recall@1 on Pitts30k-test for evaluation on the other benchmarks.
4.3 Comparisons with SOTA Methods
In this section we provide a comprehensive comparison between our proposed SAGE and a range of state-of-the-art VPR methods. The comparison includes: NetVLAD (netvlad), SFRS (sfrs), CosPlace (cosplace), MixVPR (mixvpr), EigenPlaces (eigenplaces), CricaVPR (cricavpr), SALAD (salad), BOQ (boq), SALAD-CM (salad-cm), SuperVLAD (supervlad), EMVP (emvp), and the two re-ranking (two-stage) pipelines SelaVPR (selavpr) and FoL (FoL). Re-ranking pipelines add a computationally intensive local feature matching stage. EMVP and FoL are the current SOTA for single-stage global retrieval and two-stage VPR pipelines, respectively; implementation details for compared methods are given in App. A.4.
As shown in Tab. 2 and Tab. 3, SAGE consistently outperforms previous methods across all benchmarks and evaluation metrics. In a higher-dimensional configuration (8448-d), SAGE reaches 94.5% R@1 on MSLS-val and achieves 100% R@10 on SPED, while improving R@1 by 4.3 percentage points over the previous best single-stage method (EMVP). SAGE also maintains leading performance on challenging datasets: for example, it attains 96.0% R@1 on Nordland and 83.5% R@1 on AmsterTime, representing substantial gains relative to FoL and EMVP. We also evaluate compact configurations obtained via PCA dimensionality reduction. Even under tighter budgets (2048-d and 4096-d), SAGE remains highly competitive and often matches or even surpasses recent strong baselines. For instance, the 4096-d SAGE achieves 95.6% R@1 on Pitts30k-test and 97.7% R@1 on SPED while preserving 100% R@10, demonstrating that our proposed feature amplification and epoch-wise online geo-visual sampling produce highly discriminative global descriptors. A detailed analysis of SAGE performance across varying descriptor dimensions is presented in App. A.2.
We apply t-SNE to embed spatial features from four methods into a 2D space for visual comparison. Features are extracted from $600$ images at $50$ locations. Fig. 3 displays the 2D data projections and the corresponding Average Intra-class Distance (AID), according to Equ. 8. SAGE-B has the smallest AID indicating the tightest within location clustering. More results is presented in Fig. 7.
Table 4: Comparison of parameters (M) for various VPR methods using the DINOv2-B backbone. The value in parentheses is the number of parameters in the optional cross-image encoder.
Method Total $\downarrow$ Trainable $\downarrow$ Adapter
SALAD $_{{\color[rgb]{0,0,1}\text{CVPR' 2024}}}$ 88.0 29.8
SelaVPR $_{{\color[rgb]{0,0,1}\text{ICLR' 2024}}}$ 102.8 16.2 [Uncaptioned image] 14.2
CricaVPR $_{{\color[rgb]{0,0,1}\text{CVPR' 2024}}}$ 95.7 (+11.0) 9.15 (+11.0) [Uncaptioned image] 9.2
SALAD-CM $_{{\color[rgb]{0,0,1}\text{ECCV' 2024}}}$ 88.0 29.8
SuperVLAD $_{{\color[rgb]{0,0,1}\text{NIPS' 2024}}}$ 86.6 (+11.0) 28.4 (+11.0)
EMVP $_{{\color[rgb]{0,0,1}\text{NIPS' 2024}}}$ 88.5 1.96
SAGE (Ours) 88.5 (+7.88) 1.96 (+7.88)
Tab. 4 details parameters for various VPR methods. By employing DPN for PEFT, SAGE avoids the heavy adapter modules typical of methods like SelaVPR and CricaVPR, thus achieving a markedly lower total parameters. More strikingly, since SAGE keeps the backbone frozen and exclusively fine-tunes its lightweight DPN, SoftP, and InteractHead modules, its trainable parameters is significantly smaller than that of approaches that fine-tune portions of the Transformer encoder (e.g., SALAD, SALAD-CM, and SuperVLAD). This high parameter efficiency is attained without even compromising its SOTA performance.
Fig. 5 shows qualitative results for SAGE-B and several SOTA methods in representative challenging scenarios. SAGE consistently retrieves the correct database images, while other methods often fail to capture the most discriminative cues and produce incorrect matches. Additional qualitative visualizations are provided in App. A.1, as illustrated in Fig. 8. To illustrate the comparison, we show importance heatmaps produced by SoftP, SALAD, and CFP in Fig. 5. While all three highlight prominent static landmarks, but SoftP even more effectively concentrates on subtle, fine-grained, highly discriminative regions. Additional SoftP heatmaps are provided in App. A.1, shown in Fig. 9.
Figure 3: Visualization of spatial feature clustering using t-SNE for four methods and comparison of Average Intra-class Distance (AID). Numbers next to each class indicate intra-class distance (ID).
4.4 Ablation Study
We ablate SAGE’s components using a ViT-B backbone (8448-D) and EMVP-B as the baseline (Tab. 5). On the challenging Nordland dataset, marked by severe seasonal variations, the baseline only achieves 80.8% R@1. Integrating our SoftP and OGC modules yields substantial gains, boosting R@1 to 93.6% on MSLS-val and 95.2% on Nordland. These gains suggest that SoftP enhances discriminative local feature responses, while OGC reconstructs a geo–visual graph each training epoch to expose the model to evolving hard examples aligned with the current embedding space.
Table 5: Ablation of SAGE components. All experiments use ViT-B; results reproduced in this work are marked . OGC denotes Online Graph Creation and GWS denotes Greedy Weighted Sampling.
Method Components SPED Pitts30k-test MSLS-val Nordland
Aggregation OGC GWS R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
EMVP-B CFP 91.8 96.5 97.4 93.1 96.8 97.6 93.2 96.9 97.2 80.8 90.4 93.5
SAGE-B SoftP 96.8 98.2 98.7 94.6 97.2 97.9 93.6 96.8 97.1 95.2 98.4 98.7
SoftP 96.5 97.8 98.3 93.8 96.5 97.2 92.5 96.6 96.9 94.2 97.4 97.9
CFP 97.5 98.4 98.9 94.9 97.3 98.0 93.9 97.1 97.4 95.4 98.5 98.8
SoftP 98.0 98.7 99.2 95.4 97.6 98.3 94.3 97.2 97.6 95.8 98.7 99.2
Figure 4: Qualitative results. SAGE consistently retrieves correct database images under severe challenges.
Figure 5: Visual comparison of importance heatmaps. SoftP shows a stronger focus on fine grained regions with high discriminative value than other methods overall.
Adding GWS to a SoftP configuration produces modest and unstable gains, suggesting sampling alone cannot exploit graph dynamics without Online Graph Creation. Enabling OGC and GWS while retaining CFP yields a notable improvement, for example Pitts30k-test R@1 of 94.9%, but still falls short of the configuration that includes SoftP. The full SAGE configuration achieves the best results, with R@1 of 95.4% on Pitts30k-test and R@1 of 98.0% on SPED, surpassing the baseline and intermediate variants. The greedy weighted clique expansion complements SoftP and OGC by focusing training on the most informative clusters and enhancing descriptor discriminability. As detailed in Tab. 6, we conducted an ablation study on the internal dimensions of the InteractHead to determine its optimal configuration.
Table 6: Ablation on InteractHead module. We vary the model dimension ($d_{\text{model}}$) and feed-forward dimension ($d_{\text{ff}}$).
($d_{\text{model}},d_{\text{ff}}$) SPED Pitts30k
R@1 R@5 R@10 R@1 R@5 R@10
(512, 1024) 98.8 99.5 99.7 95.5 97.6 98.2
(768, 1536) 98.6 99.3 99.7 96.0 97.9 98.4
(768, 1024) 98.9 99.7 100 95.8 97.8 98.4
To evaluate the computational cost of our dynamic sampling strategy, we conducted an ablation study comparing the runtime and performance of online versus offline graph creation. As shown in Table 7, our online approach incurs a modest 17.7% increase in per-epoch training time. However, this modest overhead is a worthwhile investment, as it translates directly to superior accuracy, with the online SAGE model achieving higher recall on both SPED and MSLS-val.
Table 7: Comparison of Online and Offline Graph Creation Strategies. Runtimes are reported per epoch for the online method. For the offline strategy, mining is a one-time cost.
Strategy Method Mining (min) Train (min) SPED MSLS-val
R@1 R@5 R@10 R@1 R@5 R@10
Online Cliquemining 4.3 25.1 90.0 95.4 96.2 94.3 96.9 97.6
SAGE (w/o GWS) 6.1 28.4 96.8 98.2 98.7 93.6 96.8 97.1
SAGE 6.2 28.4 98.9 99.7 100 94.5 97.4 97.8
Offline Cliquemining 21.6 25.1 89.5 94.9 96.1 94.2 97.2 97.4
SAGE (w/o GWS) 30.7 28.4 96.7 98.2 98.9 93.5 96.6 97.1
SAGE 30.9 28.4 98.5 99.3 99.5 94.2 97.3 97.7
Table 8: Convergence analysis on MSLS-val. SAGE’s dynamic sampling leads to superior performance in early training epochs.
Epoch SAGE (w/ CM) SAGE (Ours)
R@1 R@5 R@10 R@1 R@5 R@10
2 92.3 96.1 96.6 92.5 96.5 97.1
4 92.7 96.6 97.0 93.4 96.9 97.4
This result validates our central argument: dynamically adapting the sampling to the model’s evolving state is crucial for breaking the performance bottleneck of static mining strategies. Furthermore, the experiment highlights the efficiency of the GWS module itself, which adds negligible overhead while delivering a substantial accuracy boost. Crucially, this computational overhead is confined to the training phase. The inference process is unaffected, ensuring SAGE remains as efficient as comparable single-stage methods at deployment.
To demonstrate SAGE’s learning efficiency, we tracked its early-stage training performance on MSLS-val against a baseline using the Cliquemining (CM) strategy. As shown in Table 8, our model establishes a clear advantage by the fourth epoch (93.4% vs. 92.7% at R@1). The widening performance gap confirms that our dynamic sampling strategy fosters more efficient learning, leading to superior performance within the same epochs.
Beyond the quantitative ablations, Figure 6 presents SAGE’s qualitative failure modes. These failures typically occur under extreme conditions, such as severe viewpoint or illumination shifts, and heavy occlusion from dynamic objects.
Figure 6: Failure cases of SAGE, where the top-1 retrieval (red box) fails to match the ground truth.
5 Conclusion
We introduce SAGE, a novel framework for Visual Place Recognition that excels under drastic appearance and viewpoint variations. SAGE strategically enhances feature discriminability with a Soft Probing module and an InteractHead, while simultaneously optimizing the training process through online geo-visual graph construction and greedy weighted sampling for dynamic hard-negative mining. Extensive experiments on eight diverse benchmarks demonstrate that SAGE sets a new state-of-the-art while maintaining high parameter efficiency. We acknowledge a performance drop in highly dynamic environments. Future work will primarily focus on tackling even more extreme domain shifts and exploring the fusion of multi-modal cues like semantics, to further improve robustness.
Appendix A Appendix
A.1 Visualization of results
Figure 7: t-SNE visualization of feature clusters produced by SAGE-L and five other leading VPR methods. The features are extracted from 600 images across 50 distinct locations from the GSV-Cities dataset. Clustering compactness is quantitatively evaluated using the Average Intra-class Distance (AID), defined as the mean Euclidean distance of features to their corresponding class centroid. A lower AID signifies a more discriminative and compact feature representation for images of the same location. Notably, SAGE-L achieves the lowest AID (0.91), demonstrating its superior ability to group features from the same place.
To visually assess the quality of feature clustering, we first created a dedicated test set by selecting 600 images from 50 distinct locations (12 images per location) within the GSV-Cities dataset, ensuring coverage of diverse scenes and conditions. We then employed t-SNE to project the high-dimensional features generated by each method into a 2D space for visualization.
For a quantitative analysis of cluster compactness, we calculated the Average Intra-class Distance (AID). This metric is derived from the Intra-class Distance (ID), which measures the mean Euclidean distance of features within a class to their shared centroid $\mu_{i}$, as defined by:
$\mathrm{ID}_{i}=\frac{1}{|C_{i}|}\sum_{x\in C_{i}}\|x-\mu_{i}\|_{2},\qquad\mathrm{AID}=\frac{1}{N}\sum_{i=1}^{N}\mathrm{ID}_{i}.$ (8)
where $C_{i}$ is the set of feature vectors for location $i$. A lower AID value signifies a more compact and discriminative feature representation. As illustrated in Fig. 7, SAGE-L achieves the lowest AID, which quantitatively confirms its superior ability to generate robust features that are tightly clustered for the same location, effectively handling intra-class variations.
To highlight the practical robustness of our method, Fig. 8 presents a qualitative hcomparison between SAGE-B and seven leading VPR methods. The visualization is structured to systematically evaluate retrieval performance across six of the most common and difficult VPR challenges: severe viewpoint shifts, adverse weather, drastic lighting changes, long-term temporal differences, structural alterations, and dynamic occlusions. In these demanding scenarios, most state-of-the-art methods falter, failing to identify the correct place and retrieving visually plausible but incorrect matches (highlighted in red). In stark contrast, SAGE-B consistently retrieves the correct database image in every case (green boxes). This demonstrates the superior resilience of our approach, which stems from its ability to learn and focus on stable, truly discriminative features while effectively mitigating the impact of significant appearance variations.
The heatmap visualizations of the SoftP module, presented in Figure 9, elucidate its underlying mechanism. A clear pattern is evident across diverse scenes where the model learns to automatically suppress features from non-informative or transient sources, including the sky, road surfaces, and dynamic objects such as vehicles and pedestrians. Crucially, SoftP moves beyond concentrating on large static structures to prioritize fine-grained, stable details that offer reliable discriminative cues, for instance specific architectural features, window frames, and unique textures.
Figure 8: Qualitative comparison of SAGE-B against leading VPR methods under diverse and challenging conditions. Rows correspond to challenge category, from top to bottom: viewpoint change, weather change, lighting change, temporal change, structural change, and dynamic occlusion. Correct top-1 retrievals are indicated by a green bounding box, while incorrect matches are marked in red. The results visually confirm SAGE’s consistent and superior robustness across all scenarios.
Figure 9: Visualization of SoftP’s learned feature importance. Each pair consists of an input image and its corresponding heatmap, where warmer colors indicate regions assigned higher importance by SoftP. The visualizations demonstrate that SoftP automatically learns to focus on stable, fine-grained landmarks (e.g., building facades, structural details) while effectively ignoring non-discriminative regions (sky, road) and transient objects (vehicles, pedestrians).
A.2 Additional Results
Table 9: Performance analysis of SAGE across varying descriptor dimensions. Results on VPR benchmarks show a consistent improvement in retrieval accuracy as the dimension increases. The best and second best results for each dataset are shown in red bold and blue bold, respectively.
Method Dim SPED Pitts30k-test MSLS-val Nordland
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
SAGE (Ours) 128 84.7 92.8 95.7 88.9 94.5 95.7 81.8 91.5 93.0 56.1 72.2 78.1
256 91.6 96.7 97.5 93.0 95.9 96.7 87.2 94.6 95.7 70.5 83.6 87.8
512 94.2 98.2 99.0 94.4 96.5 97.2 91.4 95.7 96.8 80.1 90.0 92.9
1024 95.4 99.3 99.7 95.1 97.1 97.6 92.4 96.5 96.9 87.0 94.3 96.2
2048 95.6 99.2 99.7 95.4 97.4 97.9 92.6 96.9 97.7 91.2 96.6 97.8
3072 96.9 99.7 99.8 95.6 97.7 98.2 92.4 97.3 97.7 93.8 97.8 98.8
4096 97.7 99.8 100 95.6 97.7 98.3 93.7 97.3 97.8 94.4 98.2 99.0
8448 98.9 99.7 100 95.8 97.8 98.4 94.5 97.4 97.8 96.0 98.9 99.4
To further demonstrate the scalability and robustness of our SAGE framework, we present an performance analysis across a range of descriptor dimensions. This analysis spans both standard and more challenging VPR benchmarks, highlighting the consistent effectiveness of SAGE.
Tab. 9 details the performance of SAGE on four widely used VPR benchmarks (SPED, Pitts30k-test, MSLS-val, and Nordland) with descriptor dimensions varying from 128 to 8448. The results reveal a clear and consistent trend: retrieval accuracy, measured by Recall@N, systematically improves as the descriptor dimension increases. Notably, even at an intermediate dimension of 4096, SAGE achieves remarkable performance, including a perfect 100% R@10 on SPED. The results at 8448-D, such as 98.9% R@1 on SPED and 96.0% R@1 on Nordland, underscore the framework’s ability to leverage higher-dimensional feature spaces for enhanced discriminability.
Table 10: Performance of SAGE with varying descriptor dimensions on more challenging datasets.
Method Dim AmsterTime Tokyo24/7 Pitts250k‐test Eynsham
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
SAGE (Ours) 128 23.4 36.5 44.4 42.9 60.6 66.4 89.0 95.1 96.4 89.3 93.2 94.3
256 36.0 51.0 58.7 66.0 78.7 82.5 94.7 97.7 98.2 91.0 94.4 95.1
512 44.9 60.8 68.1 80.0 89.2 92.1 96.4 98.3 98.7 91.9 95.0 95.7
1024 55.6 70.0 76.9 89.5 94.6 95.9 97.3 98.9 99.2 92.4 95.5 96.2
2048 66.2 78.6 85.0 95.6 96.5 98.1 97.7 99.1 99.3 92.7 95.8 96.5
3072 73.6 85.8 89.9 94.9 98.4 98.7 98.1 99.3 99.5 92.9 96.0 96.7
4096 76.0 88.0 92.3 96.5 99.1 99.4 98.2 99.4 99.5 92.9 96.0 96.8
8448 83.5 93.3 95.4 97.5 99.1 99.4 98.4 99.4 99.7 93.1 96.2 97.0
Building on this, Tab. 10 extends the evaluation to more demanding datasets characterized by severe domain shifts, including AmsterTime (historical vs. modern), Tokyo24/7 (extreme viewpoint and time-of-day changes), Pitts250k-test (large-scale urban scenes), and Eynsham (rural route). The performance trend remains robust, with accuracy scaling gracefully with descriptor dimensionality. The improvement is particularly pronounced on AmsterTime, where the R@1 score surges from 23.4% at 128-D to 83.5% at 8448-D. This demonstrates SAGE’s exceptional capability to handle extreme appearance variations, validating the effectiveness of our proposed dynamic geo-visual graph exploration and feature enhancement strategies.
Table 11: Performance comparison on the SF-small benchmark.
Method R@1 R@5 R@10
CricaVPR $_{{\color[rgb]{0,0,1}\text{CVPR' 2024}}}$ 84.5 88.7 89.2
SALAD $_{{\color[rgb]{0,0,1}\text{CVPR' 2024}}}$ 85.7 88.2 89.7
SuperVLAD $_{{\color[rgb]{0,0,1}\text{NIPS' 2024}}}$ 85.8 89.1 89.5
SALAD-CM $_{{\color[rgb]{0,0,1}\text{ECCV' 2024}}}$ 84.0 88.0 89.8
EDTformer $_{{\color[rgb]{0,0,1}\text{TCSVT' 2025}}}$ 87.9 89.8 90.6
EMVP $_{{\color[rgb]{0,0,1}\text{NIPS' 2024}}}$ 88.2 90.6 91.1
SAGE (Ours) 89.3 91.5 91.9
To further assess the generalization capabilities of our method, we conducted experiments on the SF-small benchmark. Derived from SF-XL (cosplace), this dataset is particularly well-suited for evaluating generalization as its queries are drawn from distinct, non-adjacent locations.
Table 12: Sensitivity analysis of $\tau_{1}$ and $\tau_{2}$.
$\tau_{1}$ MSLS-val $\tau_{2}$ MSLS-val
R@1 R@5 R@10 R@1 R@5 R@10
20 93.8 97.3 97.7 $2.75\times 10^{3}$ 94.1 97.6 97.8
25 94.5 97.4 97.8 $2.88\times 10^{3}$ 94.5 97.4 97.8
30 93.7 97.0 97.4 $3.00\times 10^{3}$ 93.9 97.3 97.6
The results are presented in Table 11. SAGE achieves 89.3% at R@1, significantly outperforming all compared state-of-the-art methods. Our method utilizes two hyperparameters during Online Graph Creation: the geographic distance threshold $\tau_{1}$ and the affinity score threshold $\tau_{2}$. We conducted a sensitivity analysis on these parameters using the MSLS-val dataset, as presented in Table 12. The table also shows that the model’s performance remains stable around these optimal values, demonstrating that our method is not overly sensitive to these hyperparameters and exhibits good robustness.
To further validate the effectiveness of SAGE, we present an extended comparison against several recently published SOTA methods from top-tier venues. These include VLAD-BuFF (VLAD-BuFF), EDTFormer (jin2025edtformer), and EffoVPR (tzachor2024effovpr). The results, shown in Table 13, demonstrate that SAGE consistently outperforms these strong baselines. SAGE achieves superior performance across the majority of benchmarks. Although EffoVPR obtains a slightly higher R@1 on Tokyo24/7, it is a two-stage reranking method, which introduces additional computational overhead.
Table 13: Extended comparison with recent SOTA methods on five challenging benchmarks. Denotes two-stage reranking method.
Method SPED MSLS-val Nordland Tokyo24/7 Pitts250k-test
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
VLAD-BuFF $_{{\color[rgb]{0,0,1}\text{ECCV' 24}}}$ 91.4 95.9 96.9 91.8 96.0 96.2 85.1 93.8 96.0 96.2 98.7 99.4 95.5 98.5 99.2
EDTFormer $_{{\color[rgb]{0,0,1}\text{TCSVT' 25}}}$ 92.4 95.9 96.9 92.0 96.6 97.2 88.3 95.3 97.0 97.1 98.1 98.4 95.9 98.8 99.3
EffoVPR $_{{\color[rgb]{0,0,1}\text{ICLR' 25}}}$ 93.1 97.9 98.4 92.8 97.2 97.4 95.0 - - 98.7 98.7 98.7 - - -
SAGE (Ours) 98.9 99.7 100 94.5 97.4 97.8 96.0 98.9 99.4 97.5 99.1 99.4 98.4 99.4 99.7
A.3 Dataset Details
Pitts30k-test (pitts). The Pitts30k-test dataset is a subset of the Pittsburgh 250k dataset collected from Google Street View panoramas with GPS labels. It contains urban street-view images from Pittsburgh, Pennsylvania, USA, covering diverse city environments such as roads, bridges, and buildings, with large variations in viewpoint, season, and illumination. Each location provides multiple viewpoint images, and the dataset is primarily used for evaluating VPR models in urban scenarios.
SPED (SPED). The SPED dataset, also known as the Specific PlacEs Dataset, consists of images captured by fixed surveillance cameras over extended time periods, covering significant changes in illumination, weather, and seasons. It contains unique locations, each with hundreds of images taken at different times of day and year, providing a challenging benchmark for long-term VPR under extreme appearance variations.
MSLS-val (msls). The MSLS-val dataset is the validation split of the Mapillary Street-Level Sequences dataset, which contains street-level imagery sequences from cities worldwide captured with various devices such as smartphones, dashcams, and professional mapping cameras. The validation set covers multiple cities, seasons, and weather conditions, and is widely used for tuning and evaluating VPR models under diverse geographic and environmental variations.
Nordland (nordland). The Nordland dataset consists of front-facing video recordings from a train journey along the same railway route in Norway, captured in all four seasons: spring, summer, autumn, and winter. The route and camera viewpoints are fixed, ensuring identical spatial structure while presenting extreme appearance changes due solely to seasonal and weather differences, making it a valuable benchmark for cross-season VPR research.
Tokyo247 (tokyo247). The Tokyo247 dataset contains urban street-view images from Tokyo, Japan, primarily sourced from Google Street View panoramas with GPS labels, and is named for including both day and night imagery to represent long-term scene changes. It features dense metropolitan areas with tall buildings as well as some open spaces, offering large variations in viewpoint, illumination, and dynamic objects, and is used to evaluate VPR models in complex urban environments.
AmsterTime (amstertime). The AmsterTime dataset is a time-lapse street-view dataset of Amsterdam, containing images of the same locations captured in different years, such as 2008 and 2014, using Google Street View panoramas. It reflects long-term changes in building facades and urban infrastructure, providing a benchmark for evaluating VPR models under temporal urban transformations.
Pitts250k-test (pitts). The Pitts250k-test dataset is the test split of the Pittsburgh 250k dataset, consisting of street-view images from Pittsburgh and surrounding areas, captured from Google Street View panoramas with GPS labels. It offers diverse urban scenes with dense road networks, buildings, and bridges, and serves as a large-scale benchmark for evaluating the scalability and robustness of VPR systems.
Eynsham (eynsham; berton2022deep). The Eynsham dataset is a GPS-synchronized street-view image sequence collected along approximately 35 km of driving routes in Eynsham, Oxfordshire, UK, using a vehicle-mounted camera. It provides continuous video frames with precise ground-truth positions, making it suitable for sequence-based place recognition and loop closure detection experiments.
SF-XL (cosplace). The San Francisco eXtra Large (SF-XL) dataset is a massive and dense new benchmark designed to push visual geo-localization research towards realistic, city-wide applications. Comprising over 41 million images captured across a decade, it presents significant real-world challenges such as long-term temporal variations and a domain shift between its Street View database and crowd-sourced queries.
A.4 Compared Methods Details
NetVLAD (netvlad)https://github.com/Nanne/pytorch-NetVlad. A classic VPR method with a learnable VLAD layer pluggable into any CNN. Uses VGG-16 backbone, trained on Pitts30k, optimized via weakly supervised ranking loss, outperforming traditional methods.
SFRS (sfrs)https://github.com/yxgeee/OpenIBL. Addresses GPS noise by mining hard positives via self-supervised fine-grained region similarities, multi-generation training. Based on NetVLAD, VGG-16 backbone, trained on Pitts30k, outperforming state-of-the-art then.
CosPlace (cosplace)https://github.com/gmberton/CosPlace. Solves scalability in large-scale localization by framing training as classification. Constructs SF-XL dataset, uses CosPlace Groups. VGG-16/ResNet backbone, outputs 512D descriptors, low memory usage, suitable for city-scale applications.
MixVPR (mixvpr)https://github.com/amaralibey/MixVPR. This is a novel holistic feature aggregation method for VPR. It takes feature maps from pre-trained backbones as global features and iteratively incorporates global relationships into each feature map through stacked Feature-Mixer blocks (composed solely of multi-layer perceptrons), without the need for local or pyramidal aggregation. Using backbones like ResNet and trained on datasets such as GSV-Cities, it outperforms existing methods by a large margin with less than half the number of parameters compared to CosPlace and NetVLAD.
R2Former (r2former)https://github.com/bytedance/R2Former. This method introduces R2Former, a unified framework that employs a pure Transformer architecture to handle both global retrieval and local reranking in a single, end-to-end model. Its novel reranking module replaces slow geometric verification with a learnable Transformer that analyzes richer cues like feature correlation and attention, achieving state-of-the-art accuracy while dramatically reducing inference time and memory usage.
EigenPlaces (eigenplaces)https: //github.com/gmberton/EigenPlaces. This method embeds viewpoint robustness into learned global descriptors by clustering training data to explicitly present the model with different views of the same points of interest, without extra supervision. Using backbones like VGG-16 or ResNet with GeM pooling and trained on the SF-XL dataset, it outperforms state-of-the-art methods on most datasets, requiring $60\%$ less GPU memory for training and using $50\%$ smaller descriptors.
SelaVPR (selavpr)https://github.com/Lu-Feng/SelaVPR. It proposes a hybrid global-local adaptation method that adapts pre-trained foundation models (e.g., DINOv2) via lightweight adapters without modifying the pre-trained model parameters, efficiently generating global features for candidate retrieval and local features for re-ranking. It also introduces a mutual nearest neighbor local feature loss to avoid time-consuming spatial verification. Outperforming state-of-the-art methods on benchmarks like MSLS, it consumes only about $3\%$ of the retrieval time of RANSAC-based two-stage methods.
CricaVPR (cricavpr)https://github.com/Lu-Feng/CricaVPR. It proposes cross-image correlation-aware representation learning, using attention to correlate features of multiple images in a batch, enabling each image feature to gain useful information from others for enhanced robustness. A multi-scale convolution-enhanced adaptation method is designed to insert lightweight adapters into frozen pre-trained foundation models (e.g., DINOv2) to introduce multi-scale local information. It outperforms state-of-the-art methods by a large margin on multiple benchmarks with shorter training time.
SALAD (salad)https://github.com/serizba/salad. It reformulates NetVLAD’s soft assignment of local features to clusters as an optimal transport problem, considering both feature-to-cluster and cluster-to-feature relations, and introduces a ’dustbin’ cluster to discard non-informative features. Using DINOv2 as the backbone with fine-tuning, it trains in only 4 epochs. This single-stage method outperforms both single-stage and two-stage methods, with fast inference speed.
BoQ (boq)https://github.com/amaralibey/ Bag-of-Queries. This method learns a set of global queries and uses cross-attention to probe input features for consistent information aggregation. It supports both CNN and Vision Transformer backbones, trained on the GSV-Cities dataset. As a one-stage global retrieval method without re-ranking, surpasses two-stage methods, and is fast and efficient.
SALAD-CM (salad-cm)https://github.com/serizba/cliquemining.This work addresses the insufficient Geographic Distance Sensitivity (GDS) of existing VPR models by proposing a novel sample mining strategy. It constructs a graph of visually similar images and samples cliques (sets of geographically close images) from the graph as training batches to enhance the model’s ability to distinguish small-range geographic distances. Based on models like DINOv2 SALAD and MixVPR, trained on densely sampled datasets such as MSLS and Nordland, it significantly improves recall without increasing inference computational overhead.
SuperVLAD (supervlad)https://github.com/lu-feng/SuperVLAD. This method improves NetVLAD by removing cluster centers and using a small number of clusters , enhancing cross-domain generalization and simplifying the model. It also proposes 1-Cluster VLAD, which generates extremely low-dimensional descriptors by introducing “ghost clusters” and outperforms methods like GeM pooling with the same dimension. Using Transformer backbones (e.g., DINOv2) and trained on datasets like Pitts30k, it outperforms existing methods with lower feature dimensions.
EMVP (emvp)https://github.com/vincentqqb/EMVP. This method leverages Visual Foundation Models(e.g., DINOv2) and proposes a Centroid-Free Probing (CFP) stage that uses second-order features to better adapt VFM descriptors. It introduces a Dynamic Power Normalization (DPN) module to adaptively preserve task-specific information in both recalibration and CFP stages, forming a Parameter Efficiency Fine-Tuning (PEFT) pipeline. It achieves excellent performance on datasets like MSLS and Pitts250k, saving $64.3\%$ trainable parameters compared to existing state-of-the-art PEFT methods.
FoL (FoL)https://github.com/chenshunpeng/FoL. This two-stage VPR method models reliable discriminative regions via Extraction-Aggregation Spatial Alignment Loss (SAL) and Foreground-Background Contrast Enhancement Loss (CEL), guiding global feature generation and efficient re-ranking. It introduces a weakly supervised local feature training strategy based on pseudo-correspondences and a discriminative region-guided efficient re-ranking pipeline (xu2024local). Using DINOv2 as the backbone and trained on GSV-Cities, it outperforms existing two-stage methods on multiple benchmarks with higher computational efficiency.
VLAD-BuFF (VLAD-BuFF)https://github.com/Ahmedest61/VLAD-BuFF. This method improves visual place recognition accuracy by implementing a burst-aware weighting mechanism that discounts repetitive features to emphasize more distinctive visual cues. Simultaneously, it achieves high computational efficiency by using a pre-projection layer for local features, initialized with PCA, which enables rapid aggregation in a lower-dimensional space while maintaining high recall.
EDTFormer (jin2025edtformer)https://github.com/Tong-Jin01/EDTformer. This method introduces EDTformer, a simplified transformer decoder architecture that utilizes a set of learnable queries to efficiently decode and aggregate crucial information from image features for robust place recognition. Furthermore, it enhances the DINOv2 backbone with a novel Low-rank Parallel Adaptation (LoPA) method, which enables highly memory and parameter-efficient fine-tuning to deliver high performance with lower training costs.
EffoVPR (tzachor2024effovpr). This method introduces EffoVPR, which effectively utilizes a foundation model by extracting powerful local descriptors from its internal self-attention layers for a highly effective re-ranking process, even in a zero-shot setting. Furthermore, its single-stage approach achieves state-of-the-art performance with exceptionally compact global features by simplifying training to fine-tune only the model’s final layers, thus eliminating the need for external aggregation modules.
A.5 Limitations & Discussions & Future Work
Limitations and Discussions. Despite its strong performance, SAGE has some limitations. First, its performance can be compromised in highly dynamic scenes with significant, rapid occlusions (e.g., heavy traffic), as the model prioritizes static background features which may be temporarily obscured. Second, the Online Graph Construction introduces a computational step per epoch. While our experiments show this overhead is marginal on current benchmarks, the process could become time consuming for extremely large-scale datasets. This represents a trade-off between training efficiency and the adaptability of the sampling strategy. Third, the methodology’s effectiveness relies on the availability of reasonably accurate geographic coordinates during training. In scenarios with noisy or sparse GPS data, learning fine-grained spatial distinctions could be challenging.
Future Work. First, to better handle dynamic scenes, future work could integrate powerful foundation models like the Segment Anything Model (SAM) (SAM) to explicitly identify and mask transient objects. This would allow the model to focus purely on the stable, discriminative background context. Second, to enhance robustness against extreme domain shifts, we plan to incorporate multi-modal cues (han2025multimodal; xu20253d). Fusing visual features with semantic information from segmentation maps could provide complementary signals for more robust matching. Finally, the core principle of adaptive graph-based sampling holds promise beyond VPR. Applying this dynamic, “slow thinking” paradigm to other deep metric learning tasks, such as person re-identification or fine-grained image retrieval, could be a fruitful area of research.